OpenStudy (anonymous):

How would i find the scale factor of these numbers..Very confused need help pleaseee!!! Will give a metal...& Not asking for answer just help!!!

OpenStudy (anonymous):

58 in by 78 in 20 in by 36in 79 in by 30in 58 in by 32 in 128 in by 86 in 18 in by 12 in

OpenStudy (anonymous):

to find scale factor you gotta use relation Y=K *X where K is scale factor so lets say 78 is Y and 58 is X , plug in the values 78=K*58 K=78/58= 1.344 thats your scale factor ;) similarly you can do for the rest

OpenStudy (anonymous):

so that would be the calculations not the new dimension equation?

OpenStudy (anonymous):

exactly

OpenStudy (anonymous):

So i would divide 1.344 from each side

OpenStudy (anonymous):

for new equatioon

OpenStudy (anonymous):

yes :)

OpenStudy (anonymous):

so would be 43.1 in by 58 in

OpenStudy (anonymous):

your right :)

OpenStudy (anonymous):

Thank you:)

OpenStudy (anonymous):

your welcome ;)